Circular No. 8776

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 W2 (LONEOS) E. J. Christensen, Lunar and Planetary Laboratory, University of Ari-

zona, reports the recovery of P/2001 WF_{2} (cf. IAUC 7827) with the Catalina 0.68-m Schmidt telescope, with no sign of coma in three co-added 60-s CCD exposures on Nov. 18.3 UT or in four 30-s stacked images from Nov. 19.24–19.25. The indicated correction to the prediction on MPC 51822 is ΔT = −0.05 day.

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 18.27493  2^{h}36^{m}18.^{s}41  −8^{o}25′30′′.0  19.6
18.28226  2 36 17.41  −8 25 25.3  19.3
18.29691  2 36 15.43  −8 25 17.3
19.23653  2 34 13.36  −8 16 31.1  19.5
19.24258  2 34 12.68  −8 16 29.0  19.6
19.24869  2 34 11.82  −8 16 25.4  19.6
19.25475  2 34 10.95  −8 16 22.5  19.8

The following orbital elements linking the two apparitions are by B. G. Marsden, Smithsonian Astrophysical Observatory:

Epoch = 2002 Feb. 15.0 TT

T = 2002 Jan. 29.8481 TT  ω = 51.^{o}3524
e = 0.666722  Ω = 75.1326 }2000.0
q = 0.976351 AU  i = 16.9227

a = 2.929544 AU n^{o} = 0.1965641 P = 5.01 years Epoch = 2007 Jan. 20.0 TT

T = 2007 Feb. 6.1422 TT  ω = 51.^{o}4476
e = 0.665924  Ω = 75.0606 }2000.0
q = 0.979669 AU  i = 16.9051

a = 2.932476 AU n^{o} = 0.1962693 P = 5.02 years

SUPERNOVAE 2006nf–2006nq Eleven new type-Ia supernovae (designated 2006nf–2006np, all in the

magnitude range g = 21.0–23.0) found between Oct. 17 and Nov. 10 by the Sloan Digital Sky Survey II collaboration have been reported by J. Frieman and announced on CBET 740. The “Nearby Supernova Factory” collaboration reports the discovery of the type-II supernova 2006nq (R ∼ 19.7) in NEAT images obtained on Nov. 13.2 UT (details on CBET 741).

2006 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green